# DiffFNO: Diffusion Fourier Neural Operator

### v1

@@ 0.9 Refer to caption
11footnotetext: Corresponding author
@@ 1 Introduction
Arbitrary-scale SR models [17, 24, 7], which can upsample images at user-defined scales beyond those seen in training, have gained attention in recent years. Methods involving attention mechanisms [5] and representing images as continuous functions [10] have been explored. Operator-learning methods such as Super-Resolution Neural Operators (SRNO) [54] and HiNOTE [35] have further advanced this field. However, challenges remain due to the inherent differences between physics simulations—where operator learning excels [26]—and real-world images, computational demands, and the difficulty in restoring high-frequency details.
@@ 2 Related Work
Building upon these advancements, DiffFNO existing limitations by the Mode Rebalancing mechanism in WFNO. Instead of being truncated, all Fourier modes are preserved, with additional learnable weights to modulate their impact on reconstruction. Fine-grained feature representation is further enhanced by AttnNO, which captures local details by processing data directly in the spatial domain
Diffusion-Based SR and Efficient Sampling. Diffusion models have gained prominence as powerful generative models capable of producing high-quality images through iterative denoising techniques [14, 42]. In the context of SR, diffusion models have been employed to model the conditional distribution of HR images given LR inputs, achieving higher resolutions through the progressive enhancement of image details [41, 40].
Despite their effectiveness, diffusion models are computationally intensive due to the large number of time steps required in the reverse diffusion process. Such computational demands pose significant challenges for practical applications [20], especially in real-time or resource-constrained settings. To mitigate this issue, various acceleration techniques have been proposed.
(i) Deterministic Sampling via ODE Solvers: By reformulating the stochastic reverse diffusion as a deterministic ODE, advanced ODE solvers can be employed to reduce the number of sampling steps [21]. Methods like Denoising Diffusion Implicit Models (DDIM) [45] and DPM-Solver [33, 2] have demonstrated the ability to generate high-quality images with significantly fewer steps.
(ii) Operator Learning for Fast Sampling: Integrating neural operators into diffusion models accelerates sampling by learning the solution operator of the reverse diffusion process [59, 8].
(iii) Progressive Distillation: Training a distilled model to approximate the behavior of the full diffusion model allows faster sampling with fewer steps [43, 37]. Although effective, this method may require extensive retraining and potentially compromise image quality for increased speed.
@@
DiffFNO adopts (i) due to its simplicity and the robustness of established numerical methods. We also propose the ATS strategy, which adjusts integration step sizes adaptively to balance speed and quality.
@@ 3.2 Weighted Fourier Neural Operator
where $\mathbf{x}_{\text{HR}}(\mathbf{r})$ is the output function (e.g., the super-resolved image). The FNO models $\mathcal{G}$ by stacking layers of the form:
@@
where $\gamma_{l}$ is a learnable scalar parameter at layer $l$ that controls the strength of the weighting; $\alpha$ is a hyperparameter (or optionally, a learnable parameter) that determines how the weight scales with the frequency magnitude $\|\boldsymbol{\xi}\|$. $\mathbf{w}(\boldsymbol{\xi})$ assigns higher weights to higher frequencies when $\alpha>0$, thus emphasizing high-frequency components. This yields an updated $\mathcal{K}_{l}$:
@@ 3.3 Gated Fusion Mechanism
The fused feature map $\mathbf{v}_{\text{fused}}\in\mathbb{R}^{B\times H\times W\times C}$ is obtained by performing an element-wise weighted sum of the two feature maps:
@@
Gated Fusion Mechanism brings two advantages compared to a naive concatenation strategy: (i) Captures complementary Information: WFNO models global dependencies through spectral convolutions, effectively modeling long-range interactions and overall structure. In contrast, AttnNO excels at capturing local dependencies and fine-grained details via attention mechanisms. (ii) Balances contributions dynamically: Gated Fusion Mechanism allows the network to learn the importance of each feature map at each spatial location, dynamically balancing global and local information.
@@ 3.5 Adaptive Time-Step
  1. 1.
    Adaptive Time Step Selection Using a Learned Function. Optimizing the allocation of time steps based on data characteristics has been explored in previous works [28, 53]. We discretize the time interval $[0,T]$ into $N$ non-uniform time steps $\{t_{i}\}_{i=0}^{N}$, where $t_{0}=0$ and $t_{N}=T$. We introduce a learned function $\phi_{\psi}(t)$ as a weighted sum of polynomial basis functions, where the weightsparameterized by a set of learnable coefficients $\psi=\{\psi_{1},\psi_{2},\dots,\psi_{K}\}$, which adaptively determines the distribution of time steps based on the data characteristics.
@@
    Since $\phi_{\psi}(t)$ is monotonically increasing, its inverse function $\phi_{\psi}^{-1}(s)$ exists and can be efficiently computed using numerical methods.
  2. 2.
    Neural Operator Score Network. The score function, representing the gradient of the log probability density $\log p_{t}(\mathbf{x})$, is approximated using a neural network $s_{\theta}(\mathbf{x},t)$ parameterized by $\theta$:
@@
    In our architecture, $s_{\theta}$ consists of: (i) An encoder that extracts features from $\mathbf{x}_{\text{LR}}$); (ii) WFNO for capturing global dependencies and high-frequency details; (iii) AttnNO for modeling local dependencies and fine-grained structures; (iv) Gated Fusion Mechanism to dynamically combine features; (v) Time embedding $\mathbf{e}(t)$ incorporating the time variable $t$ into our neural network $s_{\theta}(\mathbf{x},t)$ using sinusoidal positional embeddings [50, 14], concatenating it with encoded features along the channel dimension before WFNO and AttnNO.
  3. 3.
    Efficient Solver Implementation. We solve the reverse-time stochastic differential equation (SDE) of the diffusion process, transformed into an ODE:
@@
    Using the adaptive time steps $\{t_{i}\}_{i=0}^{N}$, we discretize the ODE. We apply the Runge-Kutta 4th-order (RK4) method, as it balances computational cost and accuracy, requiring fewer steps than lower-order methods while retaining precision.
@@ 4 Experiments
Qualitative Results. Fig. 2 shows comparisons on a BSD100 image at scaling factor $\times$12, highlighting fine-grained details like animal fur and rock textures. It also includes an Urban100 image at continuous scaling factor of $\times$7.6, featuring large structures and fine local details such as reflections on the grass. SRNO and HiNOTE capture multiscale details effectively, from the animal’s body to tiny gaps between glass panels. However, DiffFNO reconstructs crisper edges with fewer artifacts, enhancing texture in animal fur pattern and reflections. WFNO captures large-scale patterns, while AttnNO and Gated Fusion Mechanism preserve intricate textures. The diffusion process with ATS ’s ODE solver further reduces artifacts deterministically.
@@
Effect of Gated Fusion Mechanism and AttnNO. The inference time trade-off from introducing these components was alleviated by Attention-based Neural Operator running in parallel to and sharing an encoder with WFNO.
Effect of ATS ODE Solver: ATS significantly reduces the number of inference steps from 1,000 to only 30. Image quality is maintained, even slightly improved. As Fig. DiffFNO: Diffusion Fourier Neural Operator(a) shows, DiffFNO leads in both PSNR and inference time.
@@ 5 Conclusion
We propose Diffusion Fourier Neural Opeartor (DiffFNO) for arbitrary-scale image super-resolution. DiffFNO contains a Weighted Fourier Neural Operator with a Mode Rebalancing mechanism to emphasize high-frequency details. It is complemented by a Attention-based Neural Operator through a Gated Fusion Mechanism that effectively adjusts the influence of global and local features. Image reconstruction is further refined by a diffusion process augmented with an Adaptive Time-Step ODE solver that dynamically allocates time steps, drastically cutting down inference time without compromising output quality. Experiments demonstrate DiffFNO’s competitiveness in both reconstruction quality and inference time across various benchmarks, establishing a new state-of-the-art.

### v2

@@ 0.9 Refer to caption
footnotetext: Corresponding author.
Work Done during the visit at Peking University.
@@ 1 Introduction
Arbitrary-scale SR models [17, 24, 7], which can upsample images at user-defined scales beyond those seen in training, have gained attention in recent years. Methods involving attention mechanisms [5] and representing images as continuous functions [10] have been explored. Operator-learning methods such as Super-Resolution Neural Operators (SRNO) [54] and HiNOTE [35] have further advanced this field. However, the inherent differences between physics simulations and real-world images introduce challenges from computational demands to the difficulty in restoring high-frequency details.
@@ 2 Related Work
The Mode Rebalancing mechanism in WFNO overcomes these limitations. Instead of being truncated, all Fourier modes are preserved, with additional learnable weights to modulate their impact on reconstruction. Fine-grained feature representation is further enhanced by AttnNO, which captures local details by processing data directly in the spatial domain
Diffusion-Based SR and Efficient Sampling. Diffusion models have gained prominence as powerful generative models capable of producing high-quality images through iterative denoising techniques [14, 42]. In the context of SR, diffusion models have been employed to model the conditional distribution of HR images given LR inputs, achieving higher resolutions after progressive enhancement [41, 40].
Despite their effectiveness, diffusion models are computationally intensive due to the large number of time steps required in the reverse diffusion process. Such computational demands pose significant challenges for practical applications [20], especially in real-time or resource-constrained settings. Current solutions include: (i) Deterministic Sampling via ODE Solvers: By reformulating the stochastic reverse diffusion as a deterministic ODE, advanced ODE solvers can be employed to reduce the number of sampling steps [21]. Methods like Denoising Diffusion Implicit Models (DDIM) [45] and DPM-Solver [33, 2] have demonstrated the ability to generate high-quality images with significantly fewer steps. (ii) Operator Learning for Fast Sampling: Neural operators accelerates sampling by learning the solution operator of the reverse diffusion process [59, 8]. (iii) Progressive Distillation: Training a distilled model to approximate the behavior of the full diffusion model allows faster sampling with fewer steps [43, 37]. Although effective, this method may require extensive retraining and potentially compromise image quality for increased speed.
@@
DiffFNO adopts (i) for its simplicity and the robustness of established numerical methods. We also strength it with the ATS strategy, which adjusts integration step sizes adaptively to balance speed and quality.
@@ 3.2 Weighted Fourier Neural Operator
where $\mathbf{x}_{\text{HR}}(\mathbf{r})$ is the output function (e.g., the super-resolved image). The FNO models $\mathcal{G}$ by stacking:
@@
where $\gamma_{l}$ is a learnable scalar parameter at layer $l$ that controls the strength of the weighting; $\alpha$ is a hyperparameter ($0.7$ in our experiments or optionally a learnable parameter) that determines how the weight scales with the frequency magnitude $\|\boldsymbol{\xi}\|$. $\mathbf{w}(\boldsymbol{\xi})$ assigns higher weights to higher frequencies when $\alpha>0$, thus emphasizing high-frequency components. This yields an updated $\mathcal{K}_{l}$:
@@ 3.3 Gated Fusion Mechanism
The fused feature map $\mathbf{v}_{\text{fused}}\in\mathbb{R}^{B\times H\times W\times C}$ is the element-wise weighted sum of the two feature maps:
@@
Gated Fusion Mechanism brings two advantages compared to a naive concatenation strategy: (i) Captures complementary Information: WFNO models global dependencies through spectral convolutions, effectively modeling long-range interactions and overall structure. In contrast, AttnNO excels at capturing local dependencies and fine-grained details via attention mechanisms. (ii) Balances contributions dynamically: Gated Fusion Mechanism elicits the importance of each feature map at each spatial location, dynamically balancing global and local information.
@@ 3.5 Adaptive Time-Step
1. Adaptive Time Step Selection Using a Learned Function. Optimizing the allocation of time steps based on data characteristics has been explored in previous works [28, 53]. We discretize the time interval $[0,T]$ into $N$ non-uniform time steps $\{t_{i}\}_{i=0}^{N}$, where $t_{0}=0$ and $t_{N}=T$. We introduce a learned function $\phi_{\psi}(t)$ as a weighted sum of polynomial basis functions, where the weight is parameterized by a set of learnable coefficients $\psi=\{\psi_{1},\psi_{2},\dots,\psi_{K}\}$, which adaptively determines the distribution of time steps based on the data characteristics.
@@
Since $\phi_{\psi}(t)$ is monotonically increasing, its inverse function $\phi_{\psi}^{-1}(s)$ exists and can be efficiently computed.
2. Neural Operator Score Network. The score function, representing the gradient of the log probability density $\log p_{t}(\mathbf{x})$, is approximated using a neural network $s_{\theta}(\mathbf{x},t)$ parameterized by $\theta$:
@@
In our architecture, $s_{\theta}$ consists of: (i) An encoder that extracts features from $\mathbf{x}_{\text{LR}}$); (ii) WFNO for capturing global dependencies and high-frequency details; (iii) AttnNO for modeling local dependencies and fine-grained structures; (iv) Gated Fusion Mechanism to dynamically combine features; (v) Time embedding $\mathbf{e}(t)$ incorporating the time variable $t$ into our neural network $s_{\theta}(\mathbf{x},t)$ using sinusoidal positional embeddings [50, 14], concatenating it and encoded features along the channel dimension.
3. Efficient Solver. We solve the reverse-time stochastic differential equation (SDE) of the diffusion process, transformed into an ODE:
@@
The adaptive time steps $\{t_{i}\}_{i=0}^{N}$ discretize the ODE. We apply the Runge-Kutta 4th-order (RK4) method, as it balances computational cost and accuracy, requiring fewer steps than lower-order methods while retaining precision.
@@ 4 Experiments
Qualitative Results. Fig. 2 compares arbitrary-scale SR methods on a BSD100 image (scaling factor of $\times$12) with fine-grained details like animal fur and rock textures and an Urban100 (continuous scaling factor of $\times$7.6) image featuring large structures and fine local details such as reflections on the grass. SRNO and HiNOTE capture multiscale details effectively, from the animal’s body to tiny gaps between glass panels. However, DiffFNO reconstructs crisper edges with fewer artifacts, enhancing texture in animal fur pattern and reflections. WFNO captures large-scale patterns, while AttnNO and Gated Fusion Mechanism preserve intricate textures. This multiscale approach followed by a diffusion process enhanced by the ATS ’s ODE solver further reduces visual artifacts.
@@
Effect of Gated Fusion Mechanism and AttnNO. The Gated Fusion Mechanism introduces minimal computational overhead. In addition, extra computational cost incurred by Attention-based Neural Operator is effectively mitigated by running it in parallel with WFNO while employing a shared encoder.
Effect of ATS ODE Solver: ATS dramatically reduces the number of inference steps from $1,000$ to just $30$, which substantially improves the inference time while delivering competitive performance. As demonstrated in Tab. 3, this acceleration not only preserves image quality but can even lead to slight improvements. In Fig. DiffFNO: Diffusion Fourier Neural Operator (a), DiffFNO outperforms existing methods in both PSNR and inference time.
@@ 5 Conclusion
We propose Diffusion Fourier Neural Opeartor (DiffFNO) for arbitrary-scale image super-resolution. DiffFNO is made of Weighted Fourier Neural Operator with a Mode Rebalancing mechanism to emphasize high-frequency details. It is complemented by a Attention-based Neural Operator through a Gated Fusion Mechanism that effectively adjusts the influence of global and local features. Image reconstruction is further refined by a diffusion process augmented with an Adaptive Time-Step ODE solver that dynamically allocates time steps, drastically cutting down inference time without compromising output quality. Experiments demonstrate DiffFNO’s competitiveness in both reconstruction quality and inference time across various benchmarks, establishing a new state-of-the-art.